### v1

# On the convergence of N-body simulations of the Solar System

Abstract
Most direct N-body integrations of planetary systems use a symplectic integrator with a fixed timestep. A large timestep is desirable because simulations run fast. However, simulations yield unphysical results if the timestep is too large. Surprisingly, no systematic convergence study has been performed on long (Gyr) timescales. In this paper we present numerical experiments to determine the minimum timestep one has to use in long-term integrations of the Solar System in order to recover the system’s fundamental secular frequencies and instability rate. We find that timesteps of up to 32 days, i.e. a third of Mercury’s orbital period, yield physical results in 5 Gyr integrations. We argue that the chaotic diffusion that drives the Solar System’s long-term evolution dominates over numerical diffusion and timestep resonances. Our results bolster confidence that most simulations in the literature are indeed converged and provide guidance on how to run time and energy efficient simulations while making sure results can be trusted.
1 Introduction
Analytical solutions and more recently numerical integrations of the Solar System, have a long history (for a historical summary of the field see e.g. Laskar, 2013). Different authors use different numerical algorithms, slightly different initial conditions, and different timesteps. Reassuringly, there is general agreement about the main conclusion: the Solar System is marginally stable. Specifically, there is a roughly 1% chance that planets will collide with each other or get ejected within 5 Gyr, approximately the time that the sun will remain on the main sequence (Laskar and Gastineau, 2009; Brown and Rein, 2020; Abbot et al., 2021).
Most modern simulations use symplectic integrators that make use of the so called Wisdom-Holman splitting (Wisdom, 1981; Wisdom and Holman, 1991; Kinoshita et al., 1991). Various implementations of the Wisdom-Holman algorithm exist, with differences in the order, the coordinate system, and other details such as whether symplectic correctors are used or not. For a review of splitting methods commonly used for planetary systems see Rein et al. (2019). What all the methods have in common is that they require the user to choose a fixed timestep at the beginning of the simulation [^1] The timestep is sometimes changed during the integration, for example when the eccentricity of a planet becomes large (Laskar and Gastineau, 2009). However, the timestep is typically not changed frequently as it can undo the advantages symplectic integrators otherwise provide.. Timesteps from 3 days up to 9 days are typically used in simulations that include the inner Solar System. Numerical convergence studies have been performed for short term integrations (Rauch and Holman, 1999; Wisdom, 2015; Boekholt and Portegies Zwart, 2015; Hernandez et al., 2022). But somewhat surprisingly a systematic convergence study has to our knowledge not been published for simulations of the full Solar System that are integrated over billions of years.
In this paper, we present such a convergence study. In contrast to other studies on symplectic integration methods, we do not rely on conserved quantities such as the conservation of energy or angular momentum to assess convergence. Whereas these quantities can give valuable hints regarding the performance of a numerical scheme, they are not interesting from an astrophysical point of view. What we do instead is look at quantities we actually care about from a physical point of view. In our case, these are a) the fundamental secular frequencies of the Solar System which are responsible for the evolution of the system over long timescales, and b) the ultimate stability of the system.
[…]
5
2 Secular Dynamics
To better interpret our numerical results, let us provide a very brief summary of the dynamics of the Solar System that has emerged in recent years (see e.g. Hoang et al., 2021, 2022; Mogavero and Laskar, 2022).
The Solar System is chaotic with a Lyapunov time of only 5 Myr (Laskar, 1989). This implies that we can only predict the precise locations of planets over short timescales, $\lesssim 100$ Myr. To make predictions about the evolution of the Solar System over longer timescales, we have to integrate an ensemble of systems in which each has slightly different initial conditions. These perturbations are typically of the order of centimetres and are thus all consistent with present day observations. Initially the differences grow exponentially on the Lyapunov timescale. Once the differences within the ensemble become macroscopic, the Solar systems enters a regime of chaotic diffusion. There are no (strong) mean-motion resonances present in the Solar System. The diffusion is primarily driven by several secular resonances (Laskar, 1990; Lithwick and Wu, 2011). As orbital elements of planets diffuse over time, in particular those of the inner Solar System, they can enter an area of parameter space where close encounters or physical collisions are possible. The probability of such a catastrophic event is about 1% (Laskar and Gastineau, 2009). When an instability occurs, we typically stop the integration.
[…]
5
3 Numerical Setup
We next describe our numerical setup. It is similar to that in other studies where the goal is to determine the Solar System’s instability rate with direct N-body integrations.
Our simulations consist of the sun and all eight planets with present day initial conditions taken from the NASA Horizons system. The initial positions of Mercury are slightly perturbed in each run between $1$cm and $1$m. We use REBOUND (Rein and Liu, 2012) and the Wisdom-Holman implementation WHFast (Rein and Tamayo, 2015), specifically the high order integrator WHCKL (Rein et al., 2019) with symplectic correctors and a modified kernel. This method has a generalized order of $O(\epsilon\,dt^{18}+\epsilon^{2}dt^{4}+\epsilon^{3}dt^{3})$ where $\epsilon\sim 10^{-3}$ for the Solar System (Wisdom et al., 1996). We include general relativistic corrections in the form of a modified solar potential (Tamayo et al., 2020). We ignore asteroids, satellites, tides, higher order gravitational harmonics, radiation effects, and stellar mass loss. We label a system as unstable when a planet gets ejected past 1000 AU or when two planets come closer than 4 times the Earth-Moon distance.
Because our goal is to perform a convergence study, we run 1200 simulations for 5 Gyr with different timesteps $dt$, logarithmically spaced between 3 and 60 days. The simulations are grouped into bins with 50 simulations each. We therefore expect to have on average 0.5 unstable simulations per bin.
Figure 1: Top panel: the fraction of simulations that went unstable within 5 Gyr. There are 50 simulations in each timestep bin. Bottom panel: the time of instability (black dots) for individual simulations. The colours indicate the three regimes we identify: converged (green), partially converged (blue), not converged (orange).
[…]
8
4 Results
[…]
5
4.1 Instability Fraction
The instability fraction of each bin as well as the instability times of individual simulations are shown in Fig. 1 as a function of the timestep. We group the simulations into three regimes. Simulations in the green regime with a timestep less than 6.5 days show the expected instability fraction of about 1%. Simulations in the blue regime with a timestep between 6.5 and 32 days have a slightly increased instability fraction of about 2%. Note that in both the green and the blue regime, instabilities happen late, after a few billion years. Simulations in the orange regime with a timestep larger than 32 days show a qualitatively and quantitatively different picture. Instabilities happen much earlier and the instability fraction is much higher.
Figure 2: The standard deviation of the secular frequency $g_{1}$ as a function of time for different timestep bins. The colours indicate the three regimes and are the same as in Fig. 1
[…]
5
4.2 Secular Frequency of Mercury
[…]
Simulations in the blue regime show a different behaviour. Within the first 150 Myr, the exponential divergence now clearly depends on the timestep. Simulations with a longer timestep diverge faster. These simulations are thus not fully converged. However, note that standard deviations of $g_{1}$ still level off to same value as those in the green bin after $\sim$300 Myr. We refer to these simulations are partially converged.
Simulations in the orange regime show once again a very different behaviour. There is no exponential divergence visible at all and the standard deviations do not approach the same values as in the simulations of the other regimes. Thus, these simulations are not converged and no longer physical.
[…]
5
4.3 Secular Frequencies after 5 Gyr
Figure 3: Cumulative distributions of the secular frequencies $g_{1}$ through $g_{5}$ in simulations of the Solar System after 5 Gyr with different timesteps. The black line corresponds to integrations using of the high order WHCKL integrator.
To assess the convergence in the distribution of fundamental frequencies we ran 80 additional simulations for each timestep bin of $3,6,12,18,24,36,48,96,$ and $192$ days. For these runs, we use the standard WHfast integrator of order $O(\epsilon dt^{2})$ without symplectic correctors. We also run one set of 80 simulations with a 6 day timestep using the WHCKL integrator. In contrast to the previous runs, we do not stop these simulations even if a close encounter or an ejection occurs. After 5 Gyr, we integrate all simulations for an additional 10 Myr with a 6 day timestep to accurately measure the fundamental frequencies with a modified Fourier transform (Šidlichovský and Nesvorný, 1996).
The secular frequencies $g_{1}$ (Mercury) through $g_{5}$ (Jupiter) at the end of the 5 Gyr integrations are shown as cumulative distributions in Fig. 3. Simulations in which an instability occurred are included in the distributions. For those the fundamental frequencies are vastly different and thus might appear outside of the bounds of the plots. The frequencies $g_{6}$, $g_{7}$, and $g_{8}$ associated with Saturn, Neptune, and Uranus are not plotted but show a very similar behaviour.
One can see that the distributions are statistically identical for simulations that use a timestep of 36 days or less. These simulations thus fall into the converged or partially converged regime. The distributions for simulations with timesteps of 48 days or longer are qualitatively different and are thus not converged. We do not observe any difference between the standard WHFast and the high order WHCKL integrator.
[…]
5
4.4 Energy Errors after 5 Gyr
We also measure the relative energy error $|\Delta E/E|$ at the end of the 5 Gyr simulations described in Sect. 4.3. The simulations with a 3 day timestep have a median energy error of $\sim 3\cdot 10^{-10}$. For larger timesteps the energy error then increases proportionally to $dt^{2}$. For simulations with a 96 day timestep it reaches $\sim 5\cdot 10^{-7}$. The median error for simulations with a timestep of 192 days is of order unity as it is dominated by the large number of unstable systems.
[…]
5
5 Discussion
Our results show that even simulations with very large timesteps – up to about a third of Mercury’s period (!) – can capture the chaotic diffusion phase of the Solar System surprisingly well (Sects. 4.2 and 4.3). The dynamical picture described in Sect. 2 allows us to explain why the instability in the partially converged (blue) regime is only slightly increased compared to the converged (green) regime (Sect. 4.1). Whereas the partially converged simulations clearly do not reproduce the initial exponential divergence and hence the Lyapunov time, they still more or less accurately capture the chaotic diffusion part which dominates over long timescales. The slightly higher instability rate in partially converged simulations can be explained by these simulations entering the chaotic diffusion part earlier. The fact that all the instabilities in the partially converged regime happen late further supports our argument that these instabilities are driven by chaotic diffusion in a very similar way to those in the fully converged regime.
Simulations which are not converged (orange) on the other hand fail to correctly capture both the initial exponential divergence and the later chaotic diffusion. The high rate of instabilities, the early instability times, and the different distribution of fundamental frequencies at late times are clear indications that a different mechanism is driving these system towards instability. Rauch and Holman (1999) showed that large timesteps can lead to so-called timestep resonances. The timescales associated with timestep resonances are significantly shorter than those associated with secular resonances, thus explaining the increased instability rate and the early onset of instabilities.
WHFast is a second order scheme and we find that the energy error indeed scales as $|\Delta E/E|\propto dt^{2}$ (Sect. 4.4). Whereas this is a confirmation that the scheme is implemented correctly, it provides no clues as to whether the physical results of the simulations are converged or not. Specifically the relative energy error does not show the qualitative differences that we observe between the converged, partially converged, and not converged regimes.
[…]
5
6 Conclusions
In this paper we presented results from several thousand 5 Gyr integrations of the Solar System with different timesteps. Our results show that simulations with timesteps of 6.5 days or less are fully converged. For fully converged simulations, no statistically significant differences can be observed even after 5 Gyr.
Surprisingly, we find that even simulations with timesteps as large as 32 days are almost indistinguishable after 5 Gyr in terms of instability fraction, time of instability, and secular frequencies. Such partially converged simulations might give unreliable results over short timescales, but the chaotic diffusion in the Solar System comes to our rescue over longer timescales. The diffusion washes out any information about the initial conditions. Any error we introduce because of the large timestep doesn’t matter as long as it remains small compared to the physical effects driving the planets’ evolution.
This argument shouldn’t come as a surprise because, regardless of the timestep we use, we always introduce some amount of numerical noise ($10^{-16}$) into the simulation at every mathematical operation due to finite floating point precision. The novel result from this work is that we were able to convincingly show that numerical artifacts several order of magnitude larger than floating point precision and originating from a finite timestep do also not affect the planets’ long-term evolution.
Previous studies (Wisdom, 2015; Hernandez et al., 2022) have argued that one might need up to 20 timesteps per pericenter passage time to avoid timestep resonances and achieve reliable results. This corresponds to a timestep of less than 3 days for the Solar System. Wisdom (2015) identifies potentially problematic timestep resonances, but looked only at relatively short term integrations (thousands of orbits) and very clean test cases (the Stark problem with one planet and a system with two giant planets). Whereas these test cases contain mean motion resonances, they do not contain the kind of secular resonances that are driving the long-term evolution of the Solar System (Mogavero and Laskar, 2022). The crucial difference is the timescale. For planets with masses similar to those of the (inner) Solar System, secular timescales are at least two orders of magnitude longer than those of mean motion resonances (Tamayo and Hadden, 2025). Thus any potential timestep resonance in the Solar System will be much weaker than in a system with strong mean motion resonances because the secular timescale and the timestep are so well separated. Another difference between such clean test cases and an integration of the full Solar System is the number of planets. In contrast to for example the Stark problem, the many different planet-planet interactions in the Solar System constantly change the orbits of planets and secular frequencies. Because of that the Solar System frequently enters a secular resonance only to shortly afterwards leave it again (Lithwick and Wu, 2011). So even if at one instance, a timestep resonance occurs, the system might not be in it long enough to be significantly affected by it. As this paper only looks at data from Solar System integrations, a primarily secularly evolving system, the results should not be blindly applied to other systems, especially those with strong mean motion resonances.
An important point to raise here is that all integrations in this paper were performed with a symplectic integrator. Because of chaotic diffusion, the Solar System is undergoing a random walk. The variations due to this random walk reach order unity on a short timescale of only a few 100 Myr (see Fig. 3). Thus there is no need to get every individual integration exactly right. Several times during a 5 Gyr integration, we loose any information about both the initial conditions and any potential numerical error we made. One only needs to make sure that numerical artifacts (e.g. timestep resonances, finite floating point precision) do not dominate over physical effects (e.g. mean-motion resonances, secular resonances, chaotic diffusion). This is further supported by the fact that the WHCKL integrations provide identical results compared to the plain WHFast integrations, despite the fact that WHCKL more accurate in short term integrations (Rein et al., 2019). We note however that it is important that any superimposed numerical effect doesn’t lead to long-term drifts. This is the reason why symplectic integrators and unbiased implementations are crucial for this task (Brouwer, 1937; Rein and Tamayo, 2015). The conclusion from this paper is that numerical artifacts are negligible for Solar System integrations if the timestep is 6.5 days or less. Even timesteps as large as 32 days only increase the instability rate by a factor of 2 over 5 Gyr because chaotic diffusion still mostly dominates over numerical artifacts.
We find that the historically often cited metric of the relative energy error can only provide a clue whether something has gone catastrophically wrong, but cannot be used to reliably determine whether physical results of simulations are converged or not. Instead, we strongly recommend that researchers always verify that the physical results they actually care about are independent of the timestep used. As we have shown in this paper, a surprisingly large timestep can result in converged results. Thus, a computationally inexpensive way to do a convergence study is to rerun a simulation with an increased timestep (which takes less time than the original simulation). If the physical results don’t change, then the simulations are likely converged. If they do change, then additional simulations with a smaller timestep should be run.
[…]
5
Software
The numerical integration methods underlying this article are part of the REBOUND package, available at https://github.com/hannorein/rebound. The code for the modified Fourier transform is available at https://boulder.swri.edu/~davidn/fmft/fmft.html. This research was made possible by the open-source projects Jupyter (Kluyver et al., 2016), iPython (Pérez and Granger, 2007), and matplotlib (Hunter, 2007; Droettboom et al., 2016).
[…]
5
Acknowledgments
We thank Daniel Tamayo, Nathan Kaib, and Sean Raymond for helpful discussions and feedback on this paper. This research has been supported by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grants RGPIN-2014-04553 and RGPIN-2020-04513. This research was enabled in part by support provided by Digital Research Alliance of Canada (formerly Compute Canada; alliancecan.ca). Computations were performed on the Niagara supercomputer (Loken et al., 2010; Ponce et al., 2019) at the SciNet HPC Consortium (scinethpc.ca). SciNet is funded by the following: the Canada Foundation for Innovation; the Government of Ontario; Ontario Research Fund – Research Excellence; and the University of Toronto.
[…]
5